Instrument Log
10 MFR          C1       CF Lamont, OK

Date

Operator

Comments

12/1/15

JEM

DQPR #4986 requests that the Logger Program be updated to v1.4. Loaded Logger Program v1.4 @2139 GMT.  RTS @2145 GMT.

11/3/15

JEM

A new program needs to be loaded onto the data logger to fix an issue with Logger ID's larger than 7999 not showing up properly.  Loaded the new program @1714 GMT.  Verified readings were all correct.  RTS @1715 GMT.

8/25/15

CAW

10M MFR due 90 day calibration checks.  Shut down 10M MFR at 1440 GMT and removed MFR head MY3227 and ran calibration checks, Reinstalled MFR head and system returned to service at 2030 GMT.

2/10/15

JEM

SDS has been unable to collect from the MFR for 9 hours.  Troubleshooting found the central cluster UPS had failed.  This caused the MOXA fiber hub to lose power.  R/R the UPS batteries.  Powered the UPS back on and communications links were restored at 1600 GMT.  RTS @1600 GMT.

6/9/14

CAW

10 Meter MFR due 90 day calibration check.  Power system down and removed MFR head for calibration. Down at 1120 GMT & back up at 1125 GMT. Completed calibration check and reinstalled head MFR head, Down at 1756 GMT and back up and RTS at 1800 GMT.  SN # MY227

2/25/14

CAW

10 meter MFR due 90 day calibration check.  Lowered tower at 1230 GMT on 2/25/14.  Removed MFR head and returned tower to its up-right position.  Completed 90 day calibration checks.  Lowered tower at 1830 GMT on 2/26/14 and installed the MFR head returned tower to normal operating position at 1840 GMT on 2/26/14.

11/18/13

CAW

10M MFR due for 90 day calibration check.  10M MFR shut down and MFR head removed for calibration check at 1625 GMT.  Calibration check completed and head reinstalled. RTS at 2245 GMT.

8/21/13

CAW

10M MFR is due 90 day calibration check.  Removed power and MFR head at 1155 GMT so it could be calibrated,  Reinstalled MFR head and returned system to operational status at 1750 GMT.

6/3/13

CAW

10 Meter MFR due 90 day calibration check.  Lowered tower at 1120 GMT and removed 10 meter MFR head for calibration.  Reinstalled MFR head and RTS at 1845 GMT.

3/29/13

JEM

The mentor requests a new program be loaded on the 10M MFR.  Loaded the new program and verified operation @1847 GMT.  The program failed to load due to multiple errors in the program.  Returned the logger to the original program and notified mentor of the problems with the new program.  RTS @1900 GMT.

3/11/13

CAM

No data had been collected from the MFR since 1318 GMT on 3/10.  Connected to datalogger with laptop and confirmed instrument was operational.  Further T/S found 10’ fiber optic cable between datalogger and junction at base of 10m tower was bad.  Replaced fiber optic cable and verified communications were restored.  RTS status O.K. at 1906 GMT.

2/27/13

CAW

10 meter MFR due 90 day calibration check.  Shut down 10 meter MFR and lowered tower on 2/27/13 at 1320 to remove the MFR head.  Put tower back up at 1325.  Completed calibration check on 10 meter MFR head and them Lowered the tower at 1635 on 2/28/13 to reinstall the MFR head, tower back up at 1640 on 2/28/13.

2/27/13

CAW

10 meter MFR due 90 day calibration check.  Shut down 10 meter MFR and lowered tower on 2/27/13 at 13:20 to remove the MFR head and pput tower back up at 13:25.  Completed calibration check on 10 meter MFR head and them Lowered the tower at 16:35 on 2/28/13 to reinstall the MFR head, tower back up at 16:40 on 2//28/13.

2/26/13

CAM

The site lost power at approximately 0350 GMT on 2/26 due to a severe winter storm.  Powered central cluster UPS off at 1730 GMT.  Power was restored at 1920 GMT.  Powered UPS on at 1945 GMT.  Instrument RTS status O.K and data collections resumed once SDS operations were restored.

2/8/13

JEM

DQPR #3616 requests the 10M IRT lens be replaced.  The tower was lowered from 1522 GMT to 1527 GMT.  10M MFR data will be effected during this time.

12/19/12

JEM

DQPR #3616 requests that the10M tower be lowered and the 10M IRT lens be cleaned.  Lowered the 10M tower @1625 GMT.  Cleaned the lens @1630 GMT.  RTS @1635 GMT.

10/31/12

JEM

The 10M IRT serial number needs to be verified.  Lowered the tower and verified the serial number as 1029 GMT.

7/26/12

CAW

10 meter MFR due for 90 day calibration check.  Removed and calibrated 10 Meter MFR Head at 1120 GMT.  Reinstalled and RTS at 1750 GMT.

5/3/12

CAW

10 meter MFR needs 90 day calibration check.  Shut down 10M MFR and removed head for calibration check and reinstalled MFR head.  System returned to service at 1805 GMT.

2/24/12

JEM

The Mentor has provided a new program for the 10M MFR.  Loaded the new program @1732 GMT.  Readings observed normal.  RTS @1745 GMT.

2/7/12

JEM

The tower needs to be lowered to measure the 10M MFR head cable length.  Dropped the tower @2055 GMT.  Measured the cable length and put the tower back up @2110 GMT.

2/7/12

JEM

While fixing a communication problem, noticed that channel 1 on the MFR head is not reading properly.  Found a loose connection in the 20 pin plug on the MFR heater controller.  Repaired the connection.  Channel 1 reading now appears normal.  RTS @1600 GMT.

2/7/12

JEM

The SDS is unable to communicate with the upgraded MFRSR.  Added a power supply to the instrument bit driver to restore communications.  Went ahead and upgraded the bit drivers to a newer variety.  Communication issue is corrected.  RTS @1530 GMT.

2/6/12

JEM

The *10M MFR is scheduled for an upgrade.  Powered the 10M MFR off at 1700 GMT.  Removed the old enclosure and all components as well as the MFR head and mounting hardware on the tower.  Installed new enclosure and components and new mounting hardware on the tower.  The tower was lowered from 2145 GMT to 2315 GMT.   Loaded the MFR program @2300 GMT.  Readings appeared normal.  RTS @2330 GMT.  SDS will have to update collections to be able to collect from a CR1000.

*Old SN’s – Head - 921  Logger - 93913  Control Box - WD27205

*New SN’s – Head - 227  CR1000 Logger - 2124  CR1000 WP - 2453  Power Supply - RA97055734.

The new control box does not have a WD sticker.

10/31/11

CAW

10M MFR is due for 90-day calibration.  Lowered tower at 1505 GMT and removed instrument for calibration.  Bad channel in head.  Sent data files to mentor for review before reinstalling.  Installed MFR at 1925 GMT on 11/1/11.

7/19/11

CAW

Instrument was removed on 7/18/11 for calibration and needs to be installed. Lowered tower at 1210 GMT, installed instrument, and raised tower.  RTS at 1225 GMT.

7/18/11

CAW

10M MFR is due for 90-day calibration.  Lowered tower at 1520 GMT and removed instrument for calibration.  Raised tower at 1525 GMT.

4/15/11

JEM

The instrument has not collected since the 0400 hour.  Due to other instruments that are down the source of the problems seems to be the Rocketport hub.  The Rocketport power light is flashing.  Power cycling the Rocketport hub did not correct the issue.  Notified the IT department of the issue.

4/5/11

CAW

MFR is due for 90-day calibration.  Lowered tower and removed MFR head at 1140 GMT.  Raised tower at 1145 GMT.  Lowered tower at 2200 GMT to reinstall MFR after it was calibrated.  Raised tower and RTS at 2207 GMT.

4/4/11

CAM

Last successful SDS collection for BRS, 10MFR and MFRC1 was at 1800 GMT on 4/3.  Power cycled the Rocket Port hub that all 3 instruments share at 1250 GMT on 4/4.  The following collection was successful at 1304 GMT.

1014/11

PAD

Last SDS collection was on 2/11/11 at 0504 GMT.  Power cycled instrument control board at 1448 GMT.  Instrument successfully collected at 1604 GMT.  RTS at 1503 GMT.

11/23/10

CAW

MFR is due for 90-day calibration.  Lowered tower and removed MFR head at 1438 GMT.  Raised tower at 1441 GMT.  Lowered tower at 2155 GMT to reinstall MFR after it was calibrated.  Raised tower and RTS at 2158 GMT.

9/7/10

JEM

A new power supply needs to be installed in the SMOS to run the PWD. The 10M MFR comms and power runs through the SMOS so 10M MFR data will be affected.  Powered off the 10M MFR @1757 GMT. Restored power to the 10M MFR @2030 GMT.

8/17/10

8/18/10

CAW

10M MFR is due for calibration.  Lowered tower and removed MFR head at 1932 GMT on 8/17.  Raised tower at 1935 GMT.  MFR was calibrated and then reinstalled on 8/18 at 1755 GMT.  RTS at 1759 GMT.

8/13/2010

JEM

The 10M MFR has not collected data since 8/12/10 at 1504 GMT.  T/S found no obvious problems with the instrument.  Replaced the data storage card on the logger board @1450 GMT.  Power cycled the instrument.  The 1504 GMT collection reinitialized the MFR.  The 1604 GMT collection was successful.

8/12/10

JES

No data has been collected from the 10M MFR since 1404 GMT today.  Checked to see if 10M MFR is getting power and it is.  Powered off at 2000 and checked power fuse and it appears to be OK.  Powered 10M MFR back on at 2002 GMT.  10M MFR initialized at 2004 GMT.  RTS @ 2010 GMT.

5/13/10

PAD

10M MFR was removed on 5/11/10, calibrated and needs to be installed.  Severe weather on 5/12/10 prevented installation.  Lowered tower and installed MFR head at 1407 GMT.  Raised tower at 1410 GMT.

5/11/10

CAW

10M MFR is due for calibration.  Lowered tower and removed MFR head at 2014 GMT to calibrate.  Raised tower at 2018 GMT.

4/5/10

PAD

Notified by facilities they were requested to take a picture of 10M MFR instrument on the tower.  Tower was lowered at 2025 GMT and raised at 2030 GMT.  RTS at 2040 GMT.

2/9/10

2/10/10

CAW

10M MFR is due for calibration.  Removed MFR head at 1852 GMT on 2/9 and started calibration.  Reinstalled after calibration was complete on 2/10 and RTS at 1652 GMT.

11/25/09

MK

Since the MFR head was reinstalled from calibration, the narrrowband and broadband measurements have flatlined. Lowered tower at 1950 GMT, removed MFR cable from head and cleaned contacts. Raised tower at 2000 and RTS at 2005 GMT.

11/10/09

PAD

Instrument was removed on 11/9/09 for calibration and has not collected since RTS at 2217 GMT.  Found power switch on control board in off position.  Powered up board.  RTS at 1512 GMT.

11/9/09

CAM

10M MFR is due for calibration.  Lowered tower at 1638 GMT and removed MFR.  Raised tower at 1639 GMT.  Lowered tower again at 2215 GMT and reinstalled MFR after it was calibrated.  Raised tower and RTS at 2217 GMT.

8/18/09

MK

10M MFR is scheduled for calibration.  Lowered tower on 8/18/09 at 1900-1905 GMT removed MFR head and started calibrations. On 8/19/09 calibrations were completed, lowered tower at 1810 GMT-1815 GMT reinstalled MFR.

6/10/09

MK

Head temp values appear to have failed again since last repair. Lowered tower at 1505 GMT, removed head cable and bent pins 16 and 18 for better connection with head connector. Reconnected, raised tower and powered back at 1520 GMT.

5/20/09

MK

Head temp is reporting low. Monitored MFR data, head temp only measuring around 2040 counts. Lowered tower at 1715 GMT, removed head connector and reconnected. Power cycled logger board and monitored data, head temp rose to 2250, suspect faulty connection on head plug.

5/12/09

JES

10M MFR is due for calibration.  Removed 10M MFR for calibration @ 1313 GMT.  Replaced 10M MFR @ 1938 GMT after it was calibrated.  RTS @ 1940 GMT.

2/6/09

MK

The lamp calibrations have been completed and the 10M MFR sensor need to be reinstalled. Lowered 10M tower at 1656 GMT, installed MFR head, raised tower at 1657 GMT and powered on MFR at 1658 GMT.

2/4/09

MK

Mentor has requested lamp calibrations for the 10M MFR. Powered off 10M MFR at 1711 GMT. Lowered tower and removed MFR head, raised tower at 1713 GMT. The MFRSR head has been taken to the  calibration facility so Craig Webb can calibrate on 2/5/09.

12/31/08

PAD

Last SDS collection was on 12/30/08 at 1804 GMT.  Instruments affected were:  mfrsr, mfr10m, and brs.  These instruments share a common rocket port.  At 1509 GMT, power cycled rocket port located at instruments and verified with computer department instruments would communicate with SDS.  RTS at 1524 GMT.

6/6/08

PAD

Last SDS collection was at 0204 GMT.  T/S found instrument has power.  Power cycled instrument control board and checked bit driver voltage.  Successful SDS collection at 1604 GMT.

12/5/07

MK

The 10M MFR is due for a standard lamp calibration. Powered off MFR at 2015 GMT. Lowered tower and removed MFR at 2017 GMT. Raised tower at 2100 GMT.

12/4/07

MK

The 10M MFR is due for a standard lamp calibration. Powered off MFR at 1830 GMT. Lowered tower and removed MFR at 1833 GMT. Raised tower at 1945 GMT.

3/20/07

PAD

Power bumps caused by area storms tripped media converters located in Optical trailer and and instrument comms box.  Following instruments were affected:  MFR10m, MFRSR, TBRG, and SIRS testbed.  Reset media converters located in Optical trailer and at instrument comm box.  RTS at 1953 GMT.

2/5/07

MK

The 10M MFR is scheduled for lamp calibration. Powered off at 1832 GMT, ran lamp calibrations and returned to service at 2019 GMT.

4/10/06

MK

The communications equipment needs to be moved and reorganized into a larger enclosure. The UPS and power supplies need to be moved and reorganized into a different enclosure. Disconnected power and communications to all instruments at E13 comms enclosure at 1525 GMT. Moved UPS and power supplies, organized all communications equipment and powered back on at 1945 GMT.

11/30/05

PAD

Last successful SDS collection was on 11/29/05 at 2103 GMT.  At 1450 GMT, found Allied Teledyne media converters located at instrument and in the optical trailer had lost communication.  Reset media converters and restored communication.   RTS at 1505 GMT.

8/17/05

MR

The communications between the CF instruments and the SDS has stopped. T/S found that the fiber optic media converter had not reset.  Reset the unit in the optical trailer and in the communications box at E13.  The communications was restored.

7/26/04

PAD

Last successful collection was on 7/23/04 at 2102 GMT.  At 1400 GMT, 10M MFR was found without power.  Storms in area on 7/23/04.  Power cycled UPS at 1408 GMT.  RTS at 1423 GMT.

4/5/04

MR

The instrument is running over the allotted time for collections.  Power cycled the instrument at 1430 GMT.  The next 3 collections where normal.  RTS at 1445 GMT.

3/10/04

PAD

DS View page shows instrument last collected data on 3/9/04 at 1602 GMT.  Power cycled datalogger at 1443 GMT.  Instrument collected data at 1602 GMT.  RTS at 1458 GMT.

12/22/03

MR

The tower needs to be lowered due to maintenance on other equipment.  Lowered the tower at 2025.  Raised the tower at 2033.  RTS at 2040 GMT.

9/17/03

KC

We're noticing an artifact in the SGP MFR10m data and the suspect is something is in the field of view of the detector. The culprit is likely the cable in front of the MFR Sensor. It's a bit difficult to interpret what is going on. We're hoping moving the cable solves the problem. Moved cable from in front of the MFR Sensor and secured the cable on the frame housing at 1621 GMT. RTS at 1636 GMT

8/19/03

MR

The instrument mentor wants the 10 M MFR data logger and head calibrated.  Lowered the tower at 1615 GMT.  Removed the MFR head and data logger.  Performed the calibration.  Reinstalled the head and data logger.   RTS at 1730 GMT.

2/4/03

JEM

The entire Central Facility has experienced a power outage at approximately 1845 GMT.  Power  was restored to the Central Facility at approximately 1945 GMT.  RTS @2000 GMT. 

11/14/02

MR

The instrument mentor wants the 10M MFR removed and a calibration test run on the instrument.  Removed the data logger and head  at 2030 GMT.  Performed a calibration test on the instrument as per mentors instructions.  Reinstalled the logger and head at 1543 GMT on 11/18/02.  E-mailed the mentor the calibration info and file.  RTS at 1555 GMT on 11/18/02.

2/19/02

MR

The instrument status page indicates that the instrument has not sent any data since 2/13/02.  T/S found that the data logger seemed locked up.  Power cycled the instrument at 1650 GMT.  The instrument came on and responded to the laptop at 1651 GMT.  SDS reinitialized the instrument at 1702 GMT.  RTS at 1705 GMT.

10/4/01

MR

Joe Michalsky wants to lower the tower again to retest and recalibrate the 10M MFRSR.   Lowered the tower at 1800 GMT.  This time the calibration indicated all but one channel on the MFRSR head was ok.  Raised the tower at 1820 GMT.  RTS at 1830 GMT.

10/4/01

MR

Joe Michalsky wants to try and calibrate the 10M MFR again today.  Lowered the tower at 1420 GMT.  Connected the LICOR calibrator and laptop up to the 10M MFR.  The results of the calibration indicate that the 10M MFR head or cable is bad and may need to be replaced.  Raised the tower at 1450 GMT.

10/3/01

MR

Joe Michalsky is on site and needs to calibrate the 10M MFR.  Lowered the tower at 2210.  Shortly after starting the calibration, the UPS shut down due to the load of the calibration equipment.  Stopped the calibration and reset the UPS.  Power was restored to the 10M MFR.  Calibration will be rerun on 10/4/01 due to time restraints.

5/4/01

MR

The 10 meter MFRSR *logger board needs to be changed out for a systems upgrade. R/R the logger board. Disabled the instrument ingest due to password change needed. RTS at 1823 GMT. *Old - 93916 *New - 93913

2/7/01

MR

The mentor called and wants the heater connections on the head to be checked out. Lowered the tower at 1600 GMT. Cleaned the connection on the MFR head and the cable. Raised the tower at 1605 GMT. RTS the system at 1615 GMT. Will notify the mentor of the actions so he can monitor the heater voltage.

1/16/01

MR

John Schmelzer called and wanted me to check the connections on the MFR for the head heater. Checked the connections and did not see anything obvious. Cleaned the connectors and reseated. Checked the head heater voltage and it was back up to normal.

11/27/00

MR

The SDS can not get any data from the data logger. Power cycled the data logger at 1730. RTS the data logger at 1745. SDS was able to collect data from the logger at 1430 GMT.

12/28/99

MK

Due to power line maintenance starting at 1500 GMT all instruments will be down unless battery power is present. At 1536 GMT the AC power was restored to all CF instruments. RTS at 1536 GMT.

9/2/99

ZW

The mentor indicated that the sensor temp varies during the day since the tower was replaced. Site Ops powered off the 10 MFR data logger at 1940 GMT. Lowered the tower at 1942 GMT. R/R the *radiometer. Site Ops raised the tower, and powered on the data logger at 2003 GMT. Status ok. RTS 2018 GMT. *Old - 916 *New - 921

6/14/99

CAM

SDS comlog reports indicated that the 10M MFR had not collected since 0711 GMT on 6/12 with error message 'Login Prompt Not Received'. Power cycled the data logger at 1345 GMT. SDS initialized the logger at 1413 GMT and normal collections resumed at 1513 GMT. RTS status O.K.

6/11/99

MW

Jim Barnard indicated that the 10M MFR head temperature was drifting during the day and he requested that the problem be investigated. The head temperature voltage at 1600 GMT was 2163mV. Checked the head temperature voltage on pin-8 of the U25 LT1079 op-amp at 1712 GMT and the voltage was 2315mV. Powered the instrument off at 1712 GMT. Installed a sensor simulator with heater element in place of the normal sensor and powered the instrument back on. The temperature voltage went to approximately 2800mV which is normal. Powered the instrument off and plugged the sensor back in. Powered the instrument back on and measured approximately 2300mV. Will drop the tower at a later time and run some tests to determine if the problem is the sensor or the cable. RTS at 1720 GMT.

5/4/99

JT

The 10 Meter tower was replaced and the 10MFR returned to service.

4/14/99

CW

The 10M tower was taken down at 2120 GMT for tower replacement/upgrade. 10M MFR was taken off line and will remain off line until tower replacement is complete.

9/9/98

SS

The 10 meter MFRSR has not collected since 9/8/98 2314 GMT. Error message is "cannot get login prompt". Checked communications with 10 Meter MFRSR from inside instrument case and the COS box. No problems found. Had the SDS operator to tip the instrument and check communications. No problems found. RTS at 1900 GMT.

8/11/98

ZW

About 1/2 of a day after the 10M MFR Data Logger was installed it stopped communication with the SDS system. At 1822 GMT Powered off the 10M MFR, and R/R the *Data Logger. Powered 10M MFR on at 1823 GMT. Status ok. RTS 1842 GMT. *Old - 93910 *New - 93916

8/6/98

GRS

Lowered the 10m tower at 1842 GMT to read some program values on the 10m IRT. Raised the tower and RTS the instrument at 1908 GMT.

8/3/98

ZW

The MFR Mentor came to calibrate the 10M MFR sensor, and Data Logger board. At 1500 GMT Site Ops lowered the tower and R/R the 10M *MFR sensor head.. Also R/R the **Data Logger. Powered on system and raised the tower at 1600 GMT. Status ok. RTS 1600 GMT. * Old - 992 *New - 916 **Old - 93916 **New - 93910

10/4/97

CAM/MDW

SDS trouble report indicates no files are being received from 10M MFR since 1731 GMT on 4/2/97 with error "Log in prompt not received". Connected directly to the MFR data logger with laptop. Received "hello" prompt and monitored MFR status O.K. RTS at 1645 GMT. SDS indicated that collections resumed at the 1800 hour. Removing and then reconnecting the comms cable seemed to have fixed the problem.

10/3/96

GRS

* R/R MFR data logger trying to correct short files. * Old - 93910 * New - 93916

9/20/96

Mentor

T/S short data files R/R E-Prom and tested. Still having problems.

8/23/96

Mentor

Installed new E-Prom and ID numbers in the MFR data logger and MF Radiometer.

3/07/96

TG/CM

1950 - 2000 GMT Installed new version of EPROM. RTS. Old version ‘J’ New version ‘L’.

2/13/96

TMG

1530 - 2002 GMT AC power secured for construction.

1/29/96

TG/CM

1432 - 2032 GMT AC power secured for construction.

1/12/96

CAM

2215 GMT Re-installed the sensor heater fuse.

1/11/96

CM/GS

1536 GMT Removed sensor heater fuse as per mentors request. T/S continuing on noisy signal problem.

12/27/95

TMG

1510 GMT AC power secured for construction.

12/19/95

MW/ZW 

Lowered tower to T/S noisy signal problem. T/S could not find problem.

11/16/95

CAM

* Installed data logger. ** Installed MFR sensor head. Tested O.K. * New - 93910 ** New - 992

7/20/95

CAM

* Removed data logger. ** Removed MFR sensor head. Left system inop. * Old - 93910 ** Old - 921